OpenStudy (anonymous):

Can someone help me with something? If Log 4 (x) = 12, then log 2 (x / 4) is equal to A. 11 B. 48 C. -12 D. 22

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@MsGerrell88

myininaya (myininaya):

solve for x then use your x you found to evaluate log_2(x/4)

OpenStudy (anonymous):

I just started learning this, can you explain it to me?

myininaya (myininaya):

\[\log_a(x)=y \text{ \implies } a^y=x\]

OpenStudy (anonymous):

\[\log_{4} x=12,4^{12=x}\] \[\log_{2} \frac{ x }{ 4 }=\log_{2} \frac{ 4^{12} }{ 4 }=\log_{2} 4^{11}=\log_{2} \left( 2^2 \right)^{11}\] i think now you can complete.

myininaya (myininaya):

so you try solving \[\log_4(x)=12\]

myininaya (myininaya):

ok nevermind

OpenStudy (anonymous):

I'm completely lost.

OpenStudy (anonymous):

Nevermind, I've got it. :) Thanks.